OpenStudy (realitywillslapyou):

Consider the functions f(x) = 3x(to the 2nd power) + 1 and g(x) = x + 3. What is the value of f[g(−1)]? A. 7 B. 37 C. 13 D. 12

OpenStudy (smokeysthename):

i THINK you have to set it up like this: 3(x+3)^2+1

OpenStudy (smokeysthename):

and if its g(-1) then you plug in -1 for x

OpenStudy (smokeysthename):

which would get you 2 so it'd be 3(2)^2+1 which is 3(4)+1 12+1=13

OpenStudy (smokeysthename):

yeah im pretty sure its c lol

OpenStudy (realitywillslapyou):

Thank you!! :D :D :D!

OpenStudy (smokeysthename):

no problem (: i just had to think about it a little

OpenStudy (realitywillslapyou):

Consider the functions f(x) = 2(x − 3) and g(x) = 3x. What is the value of (f + g)(3)? A. 0 B. 9 C. −9 D. 2

OpenStudy (smokeysthename):

give me a sec i have to work this out

OpenStudy (realitywillslapyou):

Okays

OpenStudy (smokeysthename):

ok first plug in 3 for the x values in f(x) = 2(x-3)

OpenStudy (smokeysthename):

what do get?

OpenStudy (realitywillslapyou):

B?

OpenStudy (smokeysthename):

lol you catch on quick and i didnt even have to explain the rest of it. GOOD JOB :D

OpenStudy (realitywillslapyou):

Thank you

OpenStudy (smokeysthename):

no problem (:

OpenStudy (realitywillslapyou):

Select the equations that are parallel and perpendicular to y = −3x − 1 and that pass through the point (3, 1). A. parallel: y = −3x + 6 perpendicular: y = (1/3)x + 2(2/3) B. parallel: y = −3x + 10 perpendicular: y = (1/3)x C. parallel: y = -(1/3)x perpendicular: y = −3x D. parallel: y = (1/3)x + 1 perpendicular: y = 3x − 1

OpenStudy (smokeysthename):

i dont know that one you're going to have to ask in a new question.

OpenStudy (realitywillslapyou):

Okay!

OpenStudy (realitywillslapyou):

What about this? Consider the functions f(x) = 2(x − 7) and g(x) = 2x. What is the value of (f − g)(4)? A. −6 B. 14 C. −14 D. −8